### v1

# CSPCL: Category Semantic Prior Contrastive Learning for Deformable DETR-Based Prohibited Item Detectors

Abstract
Prohibited item detection based on X-ray images is one of the most effective security inspection methods. However, the foreground-background feature coupling caused by the overlapping phenomenon specific to X-ray images makes general detectors designed for natural images perform poorly. To address this issue, we propose a Category Semantic Prior Contrastive Learning (CSPCL) mechanism, which aligns the class prototypes perceived by the classifier with the content queries to correct and supplement the missing semantic information responsible for classification, thereby enhancing the model sensitivity to foreground features. To achieve this alignment, we design a specific contrastive loss, CSP loss, which includes Intra-Class Truncated Attraction (ITA) loss and Inter-Class Adaptive Repulsion (IAR) loss, and outperforms classic N-pair loss and InfoNCE loss. Specifically, ITA loss leverages class prototypes to attract intra-class category-specific content queries while preserving necessary distinctiveness. IAR loss utilizes class prototypes to adaptively repel inter-class category-specific content queries based on the similarity between class prototypes, helping disentangle features of similar categories. CSPCL is general and can be easily integrated into Deformable DETR-based models. Extensive experiments on the PIXray and OPIXray datasets demonstrate that CSPCL significantly enhances the performance of various state-of-the-art models without increasing complexity. The code will be open source once the paper is accepted.
1 Introduction
Prohibited item detection based on X-ray images is an important security inspection measure, widely used in airports, postal services, government agencies, and border control areas. With the development of computer vision technologies, researchers have started using techniques such as image classification, object detection, and semantic segmentation to assist security personnel in examining packages imaged by security screening machines for prohibited items, thereby reducing the potential risks caused by work fatigue. However, as shown in Fig. 1, X-ray images differ from natural light images in that they exhibit unique overlapping phenomena, which cause coupling of foreground and background features. This leads to the detection results of general object detection models being easily affected by background noise.
Figure 1: Feature manifold of saw (red), knife (blue), and fireworks (green). The squares and stars represent the feature representations of prohibited items and content queries, respectively.
There are two mainstream approaches for improving the anti-overlapping detection performance of detectors in response to the overlapping phenomenon in X-ray images. Methods based on attention mechanisms, including DAM [38], RIA [33], and FCAM [13], focus on foreground information either channel-wise or spatially. Methods based on label assignment, including IAA [16], HSS [3], and LAreg[39], provide the model with more accurate candidate boxes during the training process, allowing it to concentrate on perceiving semantic information from overlapping foreground and background features. However, the aforementioned methods are all designed for CNN-based detectors and lack a decoder structure to leverage the reliable prior knowledge provided by queries to perceive specific foreground features [2, 21, 37]. Recently, researchers have found that clarifying the classification semantic information of content queries in Deformable DETR-based models can enhance the ability of the decoder to perceive foreground features in X-ray images and improve the anti-overlapping detection ability [15, 14]. AO-DETR [15] introduces the CSA label assignment strategy, which constrains content queries to detect specific categories of objects, but its portability is relatively limited. MMCL [14] proposes a plug-and-play contrastive learning mechanism that establishes sample pairs between content queries and categorizes them according to the number of classes. This method indiscriminately repels inter-class sample pairs while attracting intra-class sample pairs, thereby clarifying the semantic information of categories. However, it lacks category prior information, which may lead to content queries not aligning with the true distribution of feature manifold of categories.
As shown in Fig. 1, the features of prohibited items such as knife, saw, and fireworks exhibit distinct differences, with varying degrees of differentiation between categories. Specifically, knife and saw have more similar color, texture, and contour information compared to knife and fireworks. Consequently, the representations of knife and saw are closer together in feature space than those of knife and fireworks. Based on this observation, we propose a novel Category Semantic Prior Contrastive Learning mechanism (CSPCL), which utilizes classifier weights as specific category prototypes to provide targeted correction and guidance to the content queries responsible for each category in the decoder layers, ensuring that the content queries align with the inherent characteristic distribution of the respective prohibited items. To achieve this multi-class multi-sample alignment, we design a Category Semantic Prior (CSP) loss, which includes Intra-Class Truncated Attraction (ITA) loss and Inter-Class Adaptive Repulsion (IAR) loss, and shows superior performance compared to classic N-pair [27] loss and InfoNCE [26] loss. Specifically, we employ ITA loss to attract content queries of the same category toward their respective category prototypes, providing the queries with category semantic prior information. The ITA loss includes a gradient truncation function that stops the attraction when the similarity between the prototypes and intra-class content queries exceeds a certain threshold, thereby preventing homogenization among content queries of the same class. Similarly, the IAR loss utilizes class prototypes to repel inter-class content queries, which has a repulsion factor adaptively adjusting the repulsion strength based on the similarity between prototype pairs, thereby facilitating sufficient differentiation between content queries of similar categories.
The main contributions of our work are as follows:
  1. 1.
    We propose a plug-and-play CSPCL mechanism that uses contrastive learning to align category-specific content queries with the intrinsic feature distribution of the same category of prohibited items as perceived by the classifier, thereby enhancing the anti-overlapping detection capability of Deformable DETR-based models without increasing complexity.
  2. 2.
    We propose the CSP loss tailored for the multi-class multi-sample alignment task, which more effectively aligns class prototypes with content queries than classic contrastive losses like N-pair loss and InfoNCE loss.
  3. 3.
    The CSPCL mechanism demonstrates strong generalization across various Deformable DETR variants, such as RT-DETR [40], DINO [37], and AO-DETR [15], improving detection accuracy on two prohibited item datasets, including PIXray [23] and OPIXray [33].
2 Related Work
DETR-like Detectors. DEtection TRansformer (DETR) pioneered the removal of the Non-Maximum Suppression (NMS) post-processing step, streamlining the pipeline of object detectors. However, it still has two major drawbacks: slow training convergence and the ambiguous meaning of queries [2]. To address the issue of slow convergence, Deformable DETR [41] introduces deformable attention, which only attends to a small set of key sampling points around one selected reliable reference point, significantly improving the training convergence speed. Stale-DINO [22] uses the IoU score as ground truth to supervise the predicted classification scores of positive examples, thereby improving the stability of matching. Group-DETR [5] proposes a group-wise one-to-many assignment that conducts one-to-one assignment within each group of object queries, resulting in one ground-truth object being assigned to multiple predictions. H-DETR [9] introduces an auxiliary one-to-many matching branch during training. Co-DETR [43] integrates multiple traditional one-to-many matching strategies. To clarify the meaning of queries, Conditional DETR [24] introduces 2D reference points in the query part as reliable spatial information hints. Anchor-DETR [32] and DAB-DETR [21] introduce 4D reference boxes to further provide potential position and size information of the objects. DN-DETR [12] proposes denoising training to help the network select high-quality queries, whose spatial positions are closer to the ground truth, to predict bounding boxes. DINO [37] further proposes contrastive denoising training, teaching the network how to reject queries that are far from the ground truth.
Contrastive Learning. The core idea of contrastive learning is to train more discriminative feature representations by pulling together the representations of similar samples and pushing apart the representations of dissimilar samples. Overall, contrastive learning can be divided into unsupervised contrastive learning and supervised contrastive learning. Unsupervised contrastive learning methods can be further divided into instance-wise contrastive learning and cluster-based contrastive learning. The former is represented by methods such as InstDisc [35] and SimCLR [6], where, for any given instance, its data-augmented version is treated as an intra-class sample, while other instances are considered as inter-class samples. Typically, the InfoNCE [26] loss is used for gradient updates. The latter, represented by methods such as CC [17] and TCL [18], uses pseudo-labels generated by clustering algorithms as the basis and then employs a supervised contrastive learning framework for training. Supervised contrastive learning methods have three paradigms of loss function. The first is based on the LMMN [34], which includes Triplet loss and N-pair loss. The second is based on the Softmax function, such as AM-Softmax [31], Circle loss [28], and SupCon loss [11]. The third paradigm is based on Cross-Entropy loss, including methods like EBM [7], C2AM [36], and MMCL [14]. As far as we know, the choice among these three paradigms often depends on the specific task and the reliability of the labels, with no definitive superiority among them.
Prohibited Item Detector. Since the introduction of the first large-scale pseudo-color X-ray prohibited item detection dataset by SIXray [25], research on prohibited item object detection in X-ray images has become increasingly diverse and manifold. Prohibited item detectors are typically based on general object detectors and are improved to address the overlap phenomena in X-ray images. SIXray introduced the CHR mechanism to iteratively strip away overlapping background information. OPIXray [33] proposed the DOAM module to perceive the material and edge information of foreground objects. GADet [16] and Xdet [3] proposed the IAA and HSS label assignment strategies to mitigate the foreground-background class imbalance issue caused by overlapping phenomena. FDTNet [42] and FAPID [19] transform features into the frequency domain and then use self-attention mechanisms or convolutions to extract the texture and edge information of foreground objects. AO-DETR [15] proposed the first DETR-like model in the field of prohibited item detection, utilizing the CSA strategy to train category-specific content queries that are responsible for detecting specific types of contraband, thereby enhancing the ability to detect overlapping objects. However, the portability of CSA is limited. MMCL [14] proposed a plug-and-play contrastive learning strategy that helps clarify the category-specific semantic information of content queries in Deformable DETR-based models. However, it lacks prior knowledge of the feature distribution for each category, and the indiscriminate repulsion of inter-class content queries will lead to the misalignment between the feature distributions of queries and objects within the same category. This paper presents improvements to address this issue.
3 Proposed Method
Figure 2: Illustrating the pipeline of CSPCL plugged into one Deformable DETR [41] variant DINO [37]. (a) The overall architecture of DINO. (b) The process of updating content queries in each layer of the decoder. (c) The CSPCL mechanism uses contrastive loss to leverage classifier weights as class prototypes, thereby supplementing and refining the semantic information responsible for classification in the content queries of the decoder at layer $l$.
3.1 Explore the significance of feature alignment between objects and content queries
In this part, we examine the role of content queries in the decoder of Deformable DETR-based models and elucidate the importance of aligning object features with content queries. As shown in Fig. 2(a), using DINO as a representative model of the Deformable DETR series, given an input image, the backbone of the model first extracts multi-scale features, which are then integrated through multiple layers of encoder layers to obtain global information, commonly referred to as the encoder memory $\mathbf{M}\in\mathds{R}^{N\times C}$ [2], which is passed to the decoder. In the decoder, as depicted in Fig. 2(b), the decoder first uses the $\mathbf{M}$ to predict a set of candidate results, including classification results $\mathbf{C}\in\mathds{R}^{N\times K}$ and localization results $\mathbf{R}\in\mathds{R}^{N\times 4}$, where $K$ is the number of categories in the dataset. Then, through the query selection mechanism, the top $N_{pred}$ most reliable predictions $\mathbf{R}^{0}\in\mathds{R}^{N_{pred}\times 4}$ are filtered based on their classification confidence.
Overall, the decoder refines and updates the randomly initialized content queries $\mathbf{Q}_{c}^{0}$ through $L$ layers of decoder layers, using the encoder memory $\mathbf{M}$ and spatial prior information $\mathbf{R}^{0}$. This iterative process yields more accurate classification and localization results. The decoder layer at the $l$-th layer $\mathcal{D}^{l}$, along with the corresponding prediction head, can be expressed by the following formulas:
$\{\mathbf{Q}_{c}^{l+1},\mathbf{R}^{l+1},\mathbf{C}^{l+1}\}=\mathcal{D}^{l}(\mathbf{Q}_{c}^{l},\mathbf{R}^{l},\mathbf{M};\theta^{l}),$ (1)
$\mathbf{R}^{l+1}=\sigma(\sigma^{-1}(\mathbf{R}^{l})+\Delta\mathbf{R}^{l+1}),$ (2)
$\{\Delta\mathbf{R}^{l+1},\mathbf{C}^{l+1}\}=\mathcal{H}_{R,C}^{l}(\mathbf{Q}_{c}^{l+1};\theta^{l}_{\mathcal{H}}),$ (3)
where $l\in\{x\mid x\in\mathds{Z},0\leq x<L\}$ is the decoder block index, and $L$, topically set to 6, denotes the total number of decoder blocks. The $\theta$, $\sigma(\cdot)$, $\sigma^{-1}(\cdot)$, and $\mathcal{H}_{R,C}$ denote learnable parameters, sigmoid, inverse sigmoid [41], and prediction heads of regression and classification [2], respectively. Equation 3 shows that content queries almost directly determine the prediction results at each layer of the decoder. To analyze the update process of content queries more deeply, we further decompose the decoder in Eq. 1 into self-attention mechanisms and deformable attention mechanisms for separate explanations.
The self-attention $\mathcal{D}_{s}^{l}$ first performs element-wise addition to fuse the spatial information $\mathbf{Q}_{p}^{l}$, which is obtained by positional embedding of reference boxes $\mathbf{R}^{l}$ [37], with the content query $\mathbf{Q}_{c}^{l}$, resulting in the query of self-attention $\mathbf{Q}_{s}^{l}$ and the key of self-attention $\mathbf{K}_{s}^{l}$. This establishes a global mapping, and by extracting and aggregating features in $\mathbf{V}_{s}^{l}$ that have strong correlations with the spatial information, the output $\mathbf{Q}_{s,c}^{l}$ is obtained. The $l$-th self-attention layer can be denoted as follows:
$\mathbf{Q}_{s,c}^{l}=\mathcal{D}_{s}^{l}(\mathbf{Q}_{s}^{l},\mathbf{K}_{s}^{l},\mathbf{V}_{s}^{l};\theta_{s}^{l}),$ (4)
$\mathbf{Q}_{s}^{l}=\mathbf{K}_{s}^{l}=\mathbf{Q}_{c}^{l}+\mathbf{Q}_{p}^{l},\quad\mathbf{V}_{s}^{l}=\mathbf{Q}_{c}^{l},$ (5)
where $\mathbf{K}$ and $\mathbf{V}$ are the key and value [30], respectively.
Deformable attention $\mathcal{D}_{d}^{l}$ uses reference boxes $\mathbf{R}^{l}$ to assist the deformable queries $\mathbf{Q}_{d}^{l}$, element-wise addition result of self-attention output $\mathbf{Q}_{s,c}^{l}$ and spatial embeddings $\mathbf{Q}_{p}^{l}$, and extract feature information from the encoder memory $\mathbf{M}$ to update the content queries $\mathbf{Q}_{c}^{l+1}$. The $l$-th deformable attention layer can be denoted as follows:
$\mathbf{Q}_{c}^{l+1}=\mathbf{Q}_{d,c}^{l+1}=\mathcal{D}_{d}^{l}(\mathbf{Q}_{d}^{l},\mathbf{R}^{l},\mathbf{M};\theta_{d}^{l})$ (6)
$\mathbf{Q}_{d}^{l}=\mathbf{Q}_{s,c}^{l}+\mathbf{Q}_{p}^{l}.$ (7)
From the above process, it can be seen that content queries $\mathbf{Q}_{c}^{l}$ are continuously integrated with positional information from $\mathbf{Q}_{p}^{l}$ through self-attention and deformable attention mechanisms. As the number of iterations in the decoder layer increases, the inherent category semantics in the original content query are gradually updated and replaced by the $\mathbf{Q}_{p}^{l}$ that contains location information such as contours, edges, and shapes. This results in the loss of the content query’s role in providing classification-related semantic information, such as color and texture. Therefore, if we can supplement and correct the category semantics in the content queries, we can enhance the model’s sensitivity to the corresponding foreground object features.
3.2 Category semantic prior contrastive learning (CSPCL)
We propose the CSPCL mechanism, as illustrated in Fig. 2(c), which supplements and corrects the category semantic information of content queries by aligning them with the class prototypes perceived by the classifier weights. In the preprocessing stage, we group the content queries $\mathbf{Q}_{c}^{l}\in\mathds{R}^{N_{pred}\times M}$ from the $l$-th layer based on the number of categories $K$, resulting in $K$ groups of category-specific content queries $\mathbf{Q}_{c,k}^{l}\in\mathds{R}^{n\times M}$, where $k\in\{x\mid x\in\mathds{Z},0<x\leq K\}$, and $n$ is the quotient of the number of content queries $N_{pred}$ divided by the number of categories $K$. Next, the classifier weights $\mathbf{W}\in\mathds{R}^{K\times M}$ are dimensionally expanded and adjusted to align with the dimensions of $\mathbf{Q}_{c}^{l}$, resulting in prototypes $\mathbf{P}\in\mathds{R}^{N_{pred}\times M}$. In this way, the prototype, $\mathbf{P}_{k}\in\mathds{R}^{n\times M}$, which is responsible for category $k$, shares the same dimensions as $\mathbf{Q}_{c,k}^{l}$. The prototypes $\mathbf{P}_{k}$ and the content queries $\mathbf{Q}_{c,k}^{l}$ consist of $n$ sample pairs for each category $k$. For this multi-class multi-sample alignment task, we specifically designed the CSP loss consisting of the Intra-Class Truncated Attraction (ITA) loss and the Inter-Class Adaptive Repulsion (IAR) loss. The ITA loss attracts the prototypes to their corresponding intra-class content queries, while the IAR loss repels the prototypes from inter-class content queries. The formula for CSP loss is as follows:
$\mathcal{L}_{CSP}(\mathbf{P,Q},K)=\alpha\mathcal{L}_{ITA}(\mathbf{P,Q},K)+\beta\mathcal{L}_{IAR}(\mathbf{P,Q},K),$ (8)
wherein $\alpha$ and $\beta$ are hyperparameters used to balance two losses.
ITA loss. To attract category-specific content queries to the feature manifold of intra-class prohibited items, a simple and effective method is to compute the cross-entropy loss using the cosine similarity between category prototypes and intra-class category-specific content queries. However, even prohibited items of the same category still exhibit feature variations under different backgrounds and poses, meaning that intra-class category-specific content queries need to maintain a certain degree of variability to align with these differences, thereby improving the detection accuracy. Therefore, we specifically designed a gradient truncation mechanism to ensure that content queries retain enough variance. The overall formula for the ITA loss is as follows:
$\begin{split}\mathcal{L}_{ITA}(\mathbf{P,Q},K)=&\frac{-1}{Kn(n-1)}\sum_{k=1}^{K}\sum_{i=1}^{n}\sum_{j=1}^{n}\\ &\text{log}\Big{(}\mathcal{T}\big{(}\text{sim}(\mathbf{p}_{i}^{k},\mathbf{q}_{j}^{k}),\gamma\big{)}\Big{)},\end{split}$ (9)
$\mathcal{T}(x,\gamma)=\begin{cases}1-\gamma,&1-\gamma<x\leq 1\\ x,&0<x<1-\gamma\\ 0,&-1\leq x<0\end{cases}$ (10)
where $\text{sim}(\mathbf{p}_{i}^{k},\mathbf{q}_{j}^{k})$ is the cosine similarity between the $i$-th prototype of the $k$-th class and $j$-th content query of the $k$-th class. $\mathcal{T}(x,\gamma)$ is the gradient truncation function. As shown in Fig. 3, when the input value $x$, i.e. $\text{sim}(\mathbf{p}_{i}^{k},\mathbf{q}_{j}^{k})$, exceeds the threshold $1-\gamma$, the returned ITA loss value is constant $1-\gamma$, and its derivative becomes 00. At this point, the prototype stops attracting the intra-class content queries, allowing the content queries to retain the necessary variability and margin among queries. Therefore, by adjusting the value of $\gamma$, we can control the variance of similarities of intra-class content queries.
Figure 3: (a) and (b) are the curves of the ITA loss and its derivative, respectively.
Figure 4: The 3D surface plots of our IAR loss with different $\tau$. The samples A and B have the same $\text{sim}(\mathbf{p}_{i}^{k_{1}},\mathbf{q}_{j}^{k_{2}})$, but $\text{sim}(\mathbf{p}_{i}^{k_{1}},\mathbf{p}_{j}^{k_{2}})$ of A is smaller that of B. When $\tau=0.1$, the IAR loss values of A and B are almost the same, but when $\tau=1$, The loss value of A is much smaller than that of B.
IAR loss. Although the ITA attracts content queries toward the intra-class prototypes, some stubborn samples still drift away from the sample center, and may even appear in the feature manifold of other categories. A simple solution is to use the prototype to repel inter-class content queries indiscriminately. However, the differences between the prototypes of different categories vary naturally. As shown in Fig. 1, the feature manifolds of knife and saw have less margin than those of knife and firework, which means the former is naturally more coupled than the latter, requiring stronger repulsion by the prototype to create enough margin for clarifying the semantic information of content queries for better detection performance [1, 10, 15, 14]. Therefore, we propose the IAR loss as follows:
$\begin{split}\mathcal{L}_{IAR}(\mathbf{P,Q},K)=&\frac{-1}{K(K-1)n^{2}}\sum_{k_{1}=1}^{K}\sum_{k_{2}=1}^{K}\sum_{i=1}^{n}\sum_{j=1}^{n}\mathds{1}\left[k_{1}\neq k_{2}\right]\\ &\mathcal{R}(k_{1},k_{2})\text{log}(1-\text{sim}(\mathbf{p}_{i}^{k_{1}},\mathbf{q}_{j}^{k_{2}})),\end{split}$ (11)
$\mathcal{R}(k_{1},k_{2})=e^{1-\tau\cdot\big{(}1-\text{sim}(\mathbf{p}_{i}^{k_{1}},\mathbf{p}_{j}^{k_{2}})\big{)}},$ (12)
where $\mathds{1}[k_{1}\neq k_{2}]\in\{0,1\}$ is an indicator function. It equals $1$ if $k_{1}\neq k_{2}$, and 00 in the other case. Additionally, $\mathcal{R}(k_{1},k_{2})$ is the repulsion factor, used to adjust the strength of the repulsive effect exerted by the category prototype on content queries from different categories. The higher the similarity between the prototype of the $k_{1}$-th category $\mathbf{p}_{i}^{k_{1}}$ and the prototype of the $k_{2}$-th category $\mathbf{p}_{j}^{k_{2}}$, the stronger the repulsion exerted by the prototype $\mathbf{p}_{i}^{k_{1}}$ on the content query $\mathbf{q}_{j}^{k_{2}}$. $\tau$ is the temperature coefficient, which controls the sensitivity of the IAR loss to inter-class prototypes similarity. As shown in Fig. 4, the larger the $\tau$, the more sensitive the IAR loss becomes to differences between prototypes. This means that the prototype only repels content queries belonging to similar categories while minimizing the influence on content queries from unrelated categories, which helps align the features between prototypes and queries. When $\tau=0$, $\mathcal{R}(k_{1},k_{2})$ becomes a constant, and the prototypes indiscriminately repel content queries from all categories.
3.3 Plug CSPCL into target decoder layers
The CSPCL mechanism can be inserted into the decoder of Deformable DETR-like models in a plug-and-play manner without increasing model complexity, as described in Algorithm 1. Given the number of classes $K$ in the dataset, the content queries $\mathbf{Q}_{c}$, localization results set $R$ and classification results set $C$ of all decoder layers set $L$, the class prototypes $\mathbf{P}$, and the set of target layers $\hat{L}$ where our CSPCL mechanism will be inserted. For each decoder layer $l$, the model’s classification and localization detection results are matched with the ground truth through the Hungarian matching mechanism [2] to establish a one-to-one correspondence. Then, we compute the built-in loss $\mathcal{L}_{BASE}$ of the model for the current layer $l$ based on the matched prediction result and ground truth pairs. Additionally, we determine whether the current layer belongs to the target layers set $\hat{L}$. If it does, we use the content queries from the current layer $\mathbf{Q}_{c}^{l}$, prototypes $\mathbf{P}$, and the number of categories $K$ to compute the CSP loss according to Eq. 8. This loss is then added to the model’s inherent loss to update the current layer’s loss value $\mathcal{L}^{l}$. Finally, we use the sum of the losses from each layer, denoted as $\mathcal{L}$, to update both the model parameters and the content queries of all layers.
Algorithm 1 Plug MMCL into the Decoder.
0:    Constant $K$; Set $L$, $\hat{L}$, $\mathbf{Q}_{c}$, $\mathbf{P}$, $\mathbf{R}$, $\mathbf{C}$, $\mathbf{G}$
0:
  Initialize the total loss $\mathcal{L}$ to 0;
  for  $\forall$ decoder layer index $l\in L$  do
     $\{\mathbf{R}_{i}^{l},\mathbf{C}_{i}^{l};\mathbf{G}_{i}\}\leftarrow\mathcal{H}^{l}(\mathbf{R}^{l},\mathbf{C}^{l};\mathbf{G})$;
     $\mathcal{L}^{l}\leftarrow$ $\mathcal{L}_{BASE}(\{\mathbf{R}_{i}^{l},\mathbf{C}_{i}^{l};\mathbf{G}_{i}\})$;
     if $l\in\hat{L}$ then
        $\mathcal{L}^{l}\leftarrow\mathcal{L}^{l}+\mathcal{L}_{CSP}(\mathbf{P},\mathbf{Q}_{c}^{l},K)$;
     end if
     $\mathcal{L}\leftarrow\mathcal{L}+\mathcal{L}^{l}$;
  end for
  update model parameters and $\mathbf{Q}_{c}$ to minimize $\mathcal{L}$;
  return  $\mathbf{Q}_{c}$;
3.4 Analyze the effect of the CSPCL mechanism
Figure 5: The t-SNE visualization results of the content queries (points) and classifier weights (stars) from the first decoder layer. (a) and (b) show the visualization of content queries for DINO and DINO with the CSPCL mechanism, respectively. (c) and (d) show the visualization results of the classifier weights and content queries simultaneously reduced for the two models, DINO and DINO with the CSPCL mechanism, respectively.
Figure 6: The scatter plot and Kernel Density Estimation (KDE) joint distribution plot of prediction results of the final decoder layer. Blue and Orange represent the results of DINO and DINO with the CSPCL mechanism, respectively.
To analyze the effectiveness of the CSPCL mechanism for content queries, we visualize their t-SNE [29] dimensionality reduction distribution results with classifier weights. As shown in Fig. 5(a) and Fig. 5(c), content queries of all categories of the original DINO are scattered and evenly distributed in the feature space, while also exhibiting significant differences from the category prototypes perceived by the classifier. This suggests that during the model training process, the content queries in the decoder fail to capture the semantic information responsible for classification, which is consistent with the conclusion drawn in Sec. 3.1. In contrast, Fig. 5(b) and Fig. 5(d) demonstrate that, under the CSPCL mechanism, intra-class content queries cluster together while maintaining necessary distances, whereas inter-class content queries repel each other. Additionally, the content queries of the same category show a clear correlation with the category prototypes perceived by the model, i.e., the classifier weights. This indicates that our method successfully aligns the category-specific content queries with the inherent feature distribution of contraband, thereby addressing the deficiency in Deformable DETR-like models.
To analyze the effectiveness of the CSPCL mechanism for detection results, we visualize the IoU scores and classification scores of prediction results of DINO on the PIXray dataset, as shown in Fig. 6. We draw the scatter plot of prediction results whose classification scores and IoU scores are higher than 0.3, along with the Kernel Density Estimation (KDE) curves. Blue and orange represent the results of DINO and DINO with the CSPCL mechanism, respectively. The orange points are more concentrated and distributed further to the right compared to the blue points, indicating that under the CSPCL mechanism, the content queries capture more classification semantic prior information without losing localization features, improving the accuracy of classification results. This demonstrates that our CSPCL mechanism can help the model more effectively perceive and extract foreground information of specific classes from the overlapping features in X-ray images.
4 Experiments
4.1 Experimental setup
Datasets and evaluation metrics. We conduct extensive experiments on two large-scale publicly available X-ray prohibited item datasets, OPIXray [33] and PIXray [23], which are widely used in the field of prohibited item detection. The PIXray dataset adopts the COCO [20] evaluation metric, AP, which measures the detector’s average precision across various IoU thresholds, providing a comprehensive assessment of overall detection performance. The OPIXray dataset adopts the VOC [8] evaluation metric, mAP, which is calculated as the mean average precision across all categories at an IoU threshold of 0.5.
Implementation details. For fair, all training and testing are conducted on the same computer platform, equipped with an NVIDIA GeForce RTX 4090 GPU, an Intel Core i9-13900K CPU, 64GB of memory, running Windows 10, and using PyTorch 1.13.1. Furthermore, we utilize pre-trained ResNet-50 models from the official MMDetection [4] repository as the backbone, and all models are trained by the SGD optimizer with a learning rate of 0.01, a momentum of 0.9, and a weight decay of 0.1. In addition, all models are trained for 12 epochs, following the original training strategy, with an image size of $320\times 320$.
Table 1: Ablation study about the CSPCL mechanism on DINO [37] on the PIXray [23] dataset. $L$ means all decoder layers. The superscript ’*’ represents the optimal value of the hyper-parameter.
$\gamma$ AP AP50 AP75
$5\times 10^{-2}$ 64.9 (+0.6) 86.1 72.6
$\mathbf{5\times 10^{-3}}$ 65.4 (+1.1) 86.9 72.8
$5\times 10^{-4}$ 65.2 (+0.9) 86.6 72.6
$5\times 10^{-5}$ 64.8 (+0.5) 86.3 72.4
(a) Ablation study of ITA.
$\tau$ AP AP50 AP75
0.1 64.6 (+0.3) 85.5 72.2
0.3 65.6 (+1.3) 86.8 73.3
1 65.3 (+1.0) 86.1 73.1
3 64.5 (+0.2) 86.1 71.7
(b) Ablation study of IAR.
ITA IAR AP AP50 AP75
64.3 86.5 71.0
65.4 (+0.9) 86.9 72.8
65.6 (+1.1) 86.8 73.3
66.7 (+2.4) 87.5 74.4
(c) Ablation study of ITA and IAR results.
$\hat{L}$ AP AP50 AP75
0 66.7 (+2.4) 87.5 74.4
1 66.0 (+1.7) 87.0 73.5
5 60.6 (+2.3) 82.9 67.0
$\mathbf{L}$ 67.2 (+2.9) 88.4 74.4
(d) Ablation study of the target layer set.
$\alpha$ AP AP50 AP75
5 66.6 (+2.3) 88.1 73.7
1 67.2 (+2.9) 88.4 74.4
0.5 66.7 (+2.4) 87.5 74.4
0.1 65.1 (+0.8) 86.8 72.3
(e) $\beta=1$, $\gamma^{*}=5\times 10^{-3}$, $\tau^{*}=0.3$
$\beta$ AP AP50 AP75
5 66.3 (+2.0) 87.6 73.3
1 67.2 (+2.9) 88.4 74.4
0.5 67.5 (+3.2) 88.7 74.6
0.1 66.7 (+2.4) 87.9 73.8
(f) $\alpha^{*}=1$, $\gamma^{*}=5\times 10^{-3}$, $\tau^{*}=0.3$
4.2 Ablation experiments
To obtain the optimal hyperparameters of CSP loss, we conduct extensive ablation experiments about integrating CSPCL into DINO on the PIXray dataset. As shown in Tab. 1, we first independently insert ITA loss at the 00-th decoder layer to adjust the truncation factor $\gamma$, with the optimal truncation factor $\gamma^{*}=5\times 10^{-3}$, achieving $65.4\%$ AP on the PIXray dataset. Similarly, as shown in Tab. 1, the independent insertion of IAR loss yields the optimal temperature factor $\tau^{*}=0.3$, achieving $65.6\%$ AP. Furthermore, Tab. 1 shows that when both ITA loss and IAR loss are used simultaneously, i.e., the full version of CSP loss, the model’s AP increases by $2.4\%$, which is higher than the individual uses of ITA loss and IAR loss. This demonstrates the compatibility of the two loss functions. Subsequently, we adjust the target layer set where the CSPCL is inserted. As shown in Tab. 1, the model achieves $67.2\%$ AP by aligning the content queries of all decoder layers with the category prototypes better than aligning any single layer. Notably, when aligning only one layer, shallower layers, such as layer 00, tend to perform better. This could be because deeper layers, such as layer $5$, involve content queries that more directly contribute to detection results, and indiscriminate alignment might disrupt the spatial features perceived by the model. Finally, as shown in Tab. 1 and Tab. 1, we adjust the parameters $\alpha$ and $\beta$ in Eq. 8 to control the contribution ratio of the two losses. The optimal values for $\alpha^{*}$ and $\beta^{*}$ are $1$ and $0.5$, respectively. Overall, the CSPCL mechanism with the optimal hyperparameters can help DINO achieve the highest AP score of $67.5\%$ on PIXray.
Method AP AP50 AP75 APS APM APL
DINO 64.3 86.5 71.0 19.3 48.9 73.9
DNIO + N-pair 66.0 (+1.7) 88.1 72.2 237 51.3 75.4
DINO + InfoNCE 65.4 (+1.1) 87.4 72.3 22.1 51.1 74.6
DINO + CSP (our) 67.5 (+3.2) 88.7 74.6 23.3 52.3 76.8
Table 2: Comparing classic contrastive losses with our CSP loss.
4.3 Generalization
Contrastive loss. To verify the generalization of the CSPCL mechanism for contrastive loss, we replace the proposed CSP loss with classic contrastive losses, such as N-pair loss and InfoNCE loss, to align content queries and prototypes of the same category, and evaluate them on DINO. As shown in Tab. 2, these losses still improve DINO’s AP score on the PIXray dataset by $1.7\%$ and $1.1\%$, respectively, but are less effective than the $3.2\%$ improvement achieved by CSP loss. From these results, we draw two conclusions: (1) The strategy of aligning content queries with category prototypes in the CSPCL mechanism is generally effective for prohibited item detection in X-ray images. (2) Thanks to gradient truncation and the adaptive repulsion mechanism, CSP loss is better suited for the multi-class multi-sample aligning task compared to the aforementioned classic contrastive losses.
Models and datasets. To verify the generalization of the CSPCL mechanism for models and datasets, we apply it to multiple representative Deformable DETR variants, including the basic Deformable DETR, the powerful DINO, RT-DETR, and the contraband-specialized AO-DETR on two prohibited item datasets PIXray and OPIXray. As shown in Tab. 3, the CSPCL mechanism improves the detection accuracy of Deformable DETR, RT-DETR, DINO, and AO-DETR on the PIXray dataset by $5.2\%$, $0.4\%$, $3.2\%$, and $0.8\%$ AP, respectively. Notably, CSPCL does not impose additional burdens on model complexity such as GFLOPs and PARAMs, nor does it reduce the model inference speed (FPS). Similarly, as shown in Tab. 4, CSPCL enhances the detection accuracy of the four models on the OPIXray dataset by $3.4\%$, $0.9\%$, $0.9\%$, and $0.8\%$ mAP, respectively. These experimental results demonstrate that the CSPCL mechanism can effectively improve the Deformable DETR-based models’ ability to handle overlapping detection tasks by aligning class prototypes and content queries.
Method CSPCL AP AP50 AP75 FPS PARAMs GFLOPs
Deformable DETR 36.6 66.4 37.1 60 52.14M 13.47
41.8 (+5.2) 72.3 44.5 60 52.14M 13.47
RT-DETR 61.4 84.0 68.5 64 42.81M 17.07
61.8 (+0.4) 84.3 68.7 64 42.81M 17.07
DINO 64.3 86.5 71.0 54 58.38M 26.89
67.5 (+3.2) 88.7 74.6 54 58.38M 26.89
AO-DETR 65.6 86.1 72.0 54 58.38M 26.89
66.4 (+0.8) 86.8 73.6 54 58.38M 26.89
Table 3: Generalization analysis of CSPCL mechanism on PIXray.
Method CSPCL mAP FO ST SC UT MU
Deformable DETR 65.6 68.1 30.1 88.9 66.0 75.0
69.0 (+3.4) 78.4 38.8 86.0 62.8 79.1
RT-DETR 69.2 75.6 30.5 88.4 66.7 84.9
70.1 (+0.9) 76.0 34.4 88.6 67.4 84.3
DINO 77.0 81.7 56.3 90.0 71.4 85.5
77.9 (+0.9) 82.8 56.0 89.9 74.2 86.7
AO-DETR 77.7 82.8 57.6 89.8 71.2 86.9
78.5 (+0.8) 84.3 57.7 90.2 73.2 87.1
Table 4: Generalization analysis of CSPCL strategy on OPIXray.
4.4 Visualization
Figure 7: Visualization of deformable attention sampling points, reference points, and prediction results for corresponding group content queries in the last decoder layer. (a)–(c) represent different groups. Each sampling point is shown as a filled circle, with color indicating its attention weight, and the reference point is marked by a green cross.
To investigate the effect of the CSPCL mechanism on how content queries perceive and detect foreground prohibited items, we visualize the sampling points, reference points, and detection results in the last decoder layer for both DINO and “DINO+CSPCL” models, as shown in Fig. 7. Specifically, columns (a) and (b) demonstrate that, compared to the original DINO model, the sampling points responsible for feature extraction in the “DINO+CSPCL” model are more concentrated on the contraband objects themselves, and the confidence and localization accuracy of the detection results are higher. In column (c), under the condition of severe background overlap, the DINO model mistakenly identifies a “metal shaft” in the background as a “dart”, while the “DINO+CSPCL” focuses on the less conspicuous “lighter”. These results indicate that the CSPCL mechanism helps the model focus on foreground information and enhances its ability to the anti-overlapping detection task.
5 Conclusion
In this paper, for the prohibited item detection task based on X-ray images, we propose a plug-and-play CSPCL mechanism that aligns the classifier weights and content queries in Deformable DETR-based models. This mechanism supplements and clarifies the semantic information responsible for classification within the content queries, thereby improving the model’s ability to perceive foreground information in overlapping features. Our proposed CSP loss outperforms classic contrastive losses and is better suited for the multi-class multi-sample alignment task. The CSPCL mechanism has strong generalizability, and it can enhance the anti-overlapping detection ability of various models on two datasets by classic contrastive losses, all without increasing model complexity.